SPECTROSCOPY AND DYNAMICS OF SEMICONDUCTOR GROWTH INTERMEDIATES

\underline{DENNIS J. CLOUTHIER}, Department of Chemistry, University of Kentucky, Lexington KY 40506-0055.

Gas phase reactive intermediates play an important role in various semiconductor growth processes, such as chemical vapor deposition and plasma etching. We have been studying the spectroscopy and dynamics of such intermediates, in order to establish reliable data for experimental and theoretical attempts at optimizing industrial processes. We have succeeded in producing jet-cooled UV-visible spectra of SiF_{2}, SiCl_{2}, HSiF, HSiCl, HSiBr, GeH_{2}, GeCl_{2}, GeF_{2}, HGeCl and HGeBr with resolution sufficient to resolve the vibrational and, in most cases, the rotational structure in the S_{1} - S_{0} spectra. Analyses of these spectra have yielded the vibrational frequencies, rotational constants and sructures of the silylenes and germylenes, resolving a number of anomalies in the literature. The much weaker T_{1} - S_{0} spectra of GeF_{2}, GeCl_{2} and SiF_{2} have also been observed; in the SiF_{2} case we have obtained rotationally resolved spectra, triplet spin constants and a precise r_{0} structure. In other work relevant to SF_{6} low pressure plasma etching processes, we have detected the FS_{2} radical for the first time. The spin constants and structural parameters of the combining states were obtained from the rotational analysis of high-resolution jet spectra. Most recently, we have obtained spectra of the S_{2} - S_{0} transition of silylidene, H_{2}C=Si, and determined the ground and excited state structures and some of the vibrational frequencies. Silylidene has a variety of fascinating photophysical properties, including anomalous emission from the S_{2} state, rotational level specific quantum yields of fluorescence and extensive fluorescence quantum beats.